# Optimal and fast online change point estimation in linear regression

### v1

@@ Abstract
We consider the problem of sequential estimation of a single change point in a piecewise linear regression model under a Gaussian setup. We demonstrate that a certain CUSUM-type statistic attains the minimax optimal rates for localizing the change point. Our minimax analysis unveils an interesting phase transition from a jump (discontinuity in values) to a kink (change in slope). Specifically, for a jump, the minimax rate is of order $\log(n)/n$, whereas for a kink it scales as $\bigl{(}\log(n)/n\bigr{)}^{1/3}$, given that the sampling rate is of order $1/n$. We further introduce an algorithm for the proposed online change point detector, which requires constant computational steps and constant memory per incoming sample. Finally, the empirical performance of our method is examined on both simulated and real-world data sets. An implementation is available in the R package FLOC on GitHub.
@@ 1.1 Model and main results
In this paper, we consider the sequential detection of a change in the segmented linear regression model, where observations $X_{1},...,X_{n}$ are given by
$X_{i}\coloneqq f_{\theta}\bigl{(}\frac{i}{n}\bigr{)}+\sigma\varepsilon_{i},\qquad i=1,..,n,$ (1a)
@@
$f_{\theta}\bigl{(}\frac{i}{n}\bigr{)}\coloneqq\begin{cases}\beta_{-}\left(\frac{i}{n}-\tau\right)+\alpha_{-}&\text{if }i\leq\tau,\\ \beta_{+}\left(\frac{i}{n}-\tau\right)+\alpha_{+}&\text{if }i>\tau,\end{cases}$ (1b)
where $\theta\coloneqq\left(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+}\right)$ lies in the parameter space
$\Theta_{\delta_{0}}\coloneqq\left\{\theta=\left(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+}\right):\delta_{0}\leq\tau\leq 1-\delta_{0},\max\left(|\alpha_{+}-\alpha_{-}|,\,|\beta_{+}-\beta_{-}|\right)\geq\delta_{0}\right\},$ (1c)
for some (maybe unknown) $\delta_{0}\in\left(0,{1}/{2}\right)$. A key aspect of our analysis is the distinction between two types of structural changes in $f_{\theta}$: a jump and a kink. We classify a change in $f_{\theta}$ as a jump if $\left|\alpha_{+}-\alpha_{-}\right|\geq\delta_{0}$, indicating a (significant) discontinuity in function values, and as a kink if $\left|\alpha_{+}-\alpha_{-}\right|<\delta_{0}$ but $\left|\beta_{+}-\beta_{-}\right|\geq\delta_{0}$, indicating a (significant) discontinuity in slope without a (significant) jump in values, see Figure 1. The respective parameter spaces are denoted by $\Theta^{J}_{\delta_{0}}\coloneqq\left\{\theta\in\Theta_{\delta_{0}}:|\alpha_{+}-\alpha_{-}|\geq\delta_{0}\right\}$ for jumps and $\Theta^{K}_{\delta_{0}}\coloneqq\Theta_{\delta_{0}}\setminus\Theta^{J}_{\delta_{0}}$ for kinks.
@@
Formally, we define an online detector $\hat{\tau}$ as a Markov stopping time with respect to the filtration $\left\{\mathcal{F}_{t},1\leq t\leq n\right\}$, where $\mathcal{F}_{t}$ represents the $\sigma$-algebra generated by observations $X_{1},...,X_{t}$, and let $\mathcal{T}_{n}$ denote the set of all such online detectors. Following Korostelev and Korosteleva, (2011, Chapter 6), we aim at achieving the convergence rate of the minimax quadratic risk of detection, as the sample size $n$ goes to infinity, that is, the convergence rate to zero of
$R_{n}^{*}(\delta_{0})\;\coloneqq\;\inf_{\hat{\tau}\in\mathcal{T}_{n}}\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\right],$
which balances type I (false detection) and type II (missed detection) error as
$\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\right]=\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\mathbb{I}\left(\hat{\tau}\leq\tau\right)\right]+\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\mathbb{I}\left(\hat{\tau}>\tau\right)\right].$
This framework, rooted in in-fill asymptotics, differs slightly from conventional frameworks in sequential change point detection (see Section 1.2 below), aligning more closely with offline change point analysis (see e.g. Frick et al., 2014a, ; Niu et al.,, 2016; Truong et al.,, 2020; Cho and Kirch,, 2024), where all observations are available upfront. By framing the online problem in a manner analogous to offline analysis, our approach facilitates a rigorous comparison between online and offline setups. In addition, we put emphasis simultaneously on the statistical optimality and the computation and memory efficiency. Our contributions are two folds:
@@
    We introduce the Fast Limited-memory Optimal Change (FLOC) detector, which operates in constant time per observation and requires only constant memory, particularly independent of the sample size. The implementation of FLOC is provided in an R package available on GitHub under https://github.com/AnnikaHueselitz/FLOC.
@@
    We establish that the proposed FLOC detector achieves minimax optimal rates in quadratic risk for estimating the change point, revealing a fundamental phase transition between jump and kink scenarios. Specifically, the minimax rate scales as $\mathcal{O}\left(\log(n)/n\right)$ for a jump and as $\mathcal{O}\bigl{(}(\log(n)/n)^{1/3}\bigr{)}$ for a kink, underscoring the different levels of detectability between abrupt and more gradual structural changes.
This phase transition between a jump and a kink has been similarly observed in offline change point estimation in segmented linear models (Chen,, 2021), see also Goldenshluger et al., (2006) for general nonparametric regression models. The literature on offline change point estimation in segmented linear models is notably extensive (see e.g. Bai and Perron,, 1998, Muggeo,, 2003, Frick et al., 2014b, , and more recently Lee et al.,, 2016 and Cho et al.,, 2024 for high-dimensional extensions). In Table 1, we summarize the phase transition scenarios in segmented linear models, comparing online and offline setups in terms of optimal rates, computational efficiency and memory requirements. It is worth emphasizing that, unlike offline setups, where model validity is typically required globally over the whole domain of the function, our online detector FLOC is effective even with only local model validity around the change point. This flexibility makes our approach particularly well-suited for applications where model structures are reliable only within specific localized regions, as demonstrated with the real-world data set discussed in Section 4.3.
@@ 1.2 Related work
$\sup_{1\leq k<\infty}\mathrm{D}_{k}(\tilde{\tau})\qquad\text{subject to }\;\mathbb{E}\left[\tilde{\tau}\mid\text{no change}\right]\geq\gamma$
in the asymptotic regime of $\gamma\to\infty$. Here, $D_{k}(\cdot)$ measures the detection delay when the change occurs at $k$, and specific examples include
$\mathrm{D}_{k}(\tilde{\tau})=$ $\mathrm{ess\,sup}\,\mathbb{E}[\max(\tilde{\tau}-k+1,0)\mid X_{1},\ldots,X_{k-1};\,\text{change at }k],$ see Lorden, (1971), or (2a)
$\mathrm{D}_{k}(\tilde{\tau})=$ $\mathbb{E}[\tilde{\tau}-k+1\mid\tilde{\tau}>k;\,\text{change at }k],$ see Pollak, (1985). (2b)
Within this framework, Yao, (1993) investigated the optimal detection of a jump, with respect to (2a), in linear regression models, and Yakir et al., (1999) studied the optimal detection of a kink with respect to (2b). One might relate such results to the model (1a)–(1c) in this paper by the correspondence of $\hat{\tau}=\tilde{\tau}/n$ and $n=\gamma$.
@@
As an addition to the existing literature, the proposed online detector FLOC achieves constant computational and memory complexity per observation, while also attaining minimax optimal rates in terms of quadratic risk for segmented linear regression models simultaneously for jumps and kinks (see again Table 1).
@@ 1.3 Organization and notation
The remainder of the paper is organized as follows. In Section 2, we introduce formaly the FLOC detector and establish its convergence rates in quadratic risk of estimating the change point. The lower bounds that match the derived convergence rates are given in Section 3. Section 4 examines the empirical performance of the proposed FLOC detector on both simulated and real-world data sets. All of proofs are deferred to Section 5, and Section 6 concludes the paper with a discussion.
Throughout, we denote by $\mathbb{P}_{\tau}\left(\cdot\right)$ the probability measure and by $\mathbb{E}_{\tau}\left[\cdot\right]$ the expectation under the model (1a)–(1c) with the true change point $\tau$. In particular, $\mathbb{P}_{0}\left(\cdot\right)$ and $\mathbb{E}_{0}\left[\cdot\right]$ refer to the case with no change (i.e. $\tau=0$). For sequences $\{a_{n}\}$ and $\{b_{n}\}$ of positive numbers, we use $a_{n}=\mathcal{O}(b_{n})$ to indicate that $a_{n}\leq Cb_{n}$ for some finite constant $C>0$.
@@ 2 FLOC: Statistical theory and computational complexity
In this section, we introduce the Fast Limited-memory Optimal Change (FLOC) detector, on the basis of two test statistics, one of which is used to detect a jump and the other to detect a kink. We then derive upper bounds on the quadratic risk of FLOC as well as its computational cost.
@@ 2.1 Risk bounds
Figure 2: Simulated scenarios of a jump (upper panel) and a kink (lower panel). The black lines denote the underlying signal $f_{\theta}$ with a change at $\tau=516/n$ and the blue dots the observations drawn under a Gaussian distribution with the signal as the mean, see (1a) and (1b). The red step functions show the values of the test statistic from FLOC. The vertical dotted lines show the observations where FLOC discards a bin; the bin sizes are $N_{J}=N_{k}=2$. The pre-change signal has been estimated by the first $k=500$ observations. FLOC flags a change if the test statistic exceeds the threshold, with $\rho_{J}=1.5$ and $\rho_{K}=0.3$, displayed by the dashed lines in the plot. For a jump this happens at observation $521$ and for a kink at observation $521$ as well. For visibility, the values of threshold and test statistic are multiplied by 5 for the kink case (lower panel).
@@
The idea of the online detection algorithm is to estimate the pre-change signal $f_{-}$ at the beginning of the data sequence and subsequently to verify over the most recent time interval whether the data continues to adhere to the same underlying function. The detector is built upon certain test statistics, which report a change if the test statistics exceed pre-specified thresholds. For a jump we use the mean of the data on the most recent time interval as the test statistic, and for a kink we employ a weighted mean, tuned to estimate the slope of a linear function, with higher weights on more recent observations.
@@
Consider the data sequence $X_{1},X_{2},...,X_{n}$, and let $k,N_{J},N_{K}\leq n$ be some natural numbers depending on $n$. A minimax rate optimal choice of these parameters will be specified in Theorem 1 below; The practical choice will be discussed later in Section 4.1. The function $f_{-}$ is estimated using the first $k$ data points via least squares as follows
@@
$\hat{f}_{-}(\frac{i}{n})\coloneqq\hat{\alpha}+\hat{\beta}\frac{i}{n},\qquad\text{ with }\quad\bigl{(}\hat{\alpha},\hat{\beta}\bigr{)}\coloneqq\underset{\left(\alpha,\beta\right)\in\mathbb{R}^{2}}{\arg\min}\sum_{i=1}^{k}\left(\alpha+\beta\frac{i}{n}-X_{i}\right)^{2}.$ (3a)
We define the CUSUM jump test statistic with bin size $N_{J}$ as
$J_{m}\coloneqq\left|\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(X_{m-M_{J}+i}-\hat{f}_{-}\left(\frac{m-M_{J}+i}{n}\right)\right)\right|,\qquad k<m\leq n,$ (3b)
@@
where $M_{J}=2N_{J}+\left(m\mod N_{J}\right)$. Similarly, the CUSUM kink test statistic with bin size $N_{K}$ is defined as
$K_{m}\coloneqq\left|\frac{6}{M_{K}\left(M_{K}+1\right)\left(2M_{K}+1\right)}\sum_{i=1}^{M_{K}}i\left(X_{m-M_{K}+i}-\hat{f}_{-}\left(\frac{m-M_{K}+i}{n}\right)\right)\right|,\qquad k<m\leq n,$ (3c)
with $M_{K}=2N_{K}+\left(m\mod N_{K}\right)$. We introduce for some fixed thresholds $\rho_{J},\rho_{K}>0$ the corresponding detectors $\hat{\tau}_{J,n}$ and $\hat{\tau}_{K,n}$ as the first time the test statistics exceed their respective thresholds, more precisely,
$\hat{\tau}_{J,n}\coloneqq\begin{cases}1&\text{if }J_{m}<\rho_{J}\text{ for all}m=k+1,...,n\\ \min\left(m:J_{m}\geq t,\ k<m\leq n\right)/n&\text{else,}\end{cases}$ (3d)
@@
$\hat{\tau}_{K,n}\coloneqq\begin{cases}1&\text{if }K_{m}<\rho_{K}\text{ for all}m=k+1,...,n\\ \min\left(m:K_{m}\geq t,\ k<m\leq n\right)/n&\text{else,}\end{cases}$ (3e)
@@
$\hat{\tau}_{n}\coloneqq\min\left(\hat{\tau}_{J,n},\hat{\tau}_{K,n}\right),$ (3f)
see Figure 2 for an illutration. It is straightforward to verify that the detectors $\hat{\tau}_{J,n}$ and $\hat{\tau}_{K,n}$, as defined in (3d) and (3e), are Markov stopping times. As a consequence, FLOC, being the minimum of two Markov stopping times, is also a Markov stopping time, that is, FLOC is indeed an online detector (see also the proof in Section 5).
@@
    Let $\hat{\tau}_{J,n}$ be the CUSUM jump detector in (3b) and (3d) with bin size $N_{J}=10^{3}\log(n)/(2c^{2})$ and threshold $\rho_{J}=4c/5$. Then it holds for sufficiently large $n$
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\right]\;\leq\;r^{*}_{J},$
@@
    Let $\hat{\tau}_{K,n}$ be the CUSUM kink detector in (3c) and (3e) with bin size $N_{K}=(300/c^{2})^{1/3}n^{2/3}\log^{1/3}(n)$ and threshold $\rho_{K}=4c/5n$. Then it holds for sufficiently large $n$
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\hat{\tau}_{K,n}-\tau\right)\right)^{2}\right]\leq r^{*}_{K},$
@@
    For the detector FLOC defined in (3b)–(3f), both of the above rates hold for sufficiently large $n$.

### v2

# Online jump and kink detection in segmented linear regression: Statistical optimality meets computational efficiency

@@ Abstract
We consider the problem of sequential (online) estimation of a single change point in a piecewise linear regression model under a Gaussian setup. We demonstrate that certain CUSUM-type statistics attain the minimax optimal rates for localizing the change point. Our minimax analysis unveils an interesting phase transition from a jump (discontinuity in function values) to a kink (a change in slope). Specifically, for a jump, the minimax rate is of order $\log(n)/n$ , whereas for a kink it scales as $\bigl{(}\log(n)/n\bigr{)}^{1/3}$, given that the sampling rate is of order $1/n$. We further introduce an online algorithm based on these detectors, which optimally identifies both a jump and a kink, and is able to distinguish between them. Notably, the algorithm operates with constant computational complexity and requires only constant memory per incoming sample. Finally, we evaluate the empirical performance of our method on both simulated and real-world data sets. An implementation is available in the R package FLOC on GitHub.
@@ 1.1 Model and main results
In this paper, we consider the sequential detection of a change in the segmented linear regression model, where observations $X_{1},\ldots,X_{n}$ are given by
$X_{i}\coloneqq f_{\theta}\bigl{(}\frac{i}{n}\bigr{)}+\sigma\varepsilon_{i},\qquad i=1,\ldots,n,$ (1a)
@@
$f_{\theta}\bigl{(}\frac{i}{n}\bigr{)}\coloneqq\begin{cases}\beta_{-}\left(\frac{i}{n}-\tau\right)+\alpha_{-}&\text{if }\frac{i}{n}\leq\tau,\\ \beta_{+}\left(\frac{i}{n}-\tau\right)+\alpha_{+}&\text{if }\frac{i}{n}>\tau,\end{cases}$ (1b)
where $\theta=\left(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+}\right)$ lies in the parameter space
$\Theta_{\delta_{0}}\coloneqq\bigl{\{}\theta=\left(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+}\right):\delta_{0}\leq\tau\leq 1-\delta_{0},\max\left(|\alpha_{+}-\alpha_{-}|,\,|\beta_{+}-\beta_{-}|\right)\geq\delta_{0}\bigr{\}},$ (1c)
for some (maybe unknown) $\delta_{0}\in\left(0,{1}/{2}\right)$. A key aspect of our analysis is the distinction between two types of structural changes in $f_{\theta}$: a jump and a kink. We classify a change in $f_{\theta}$ as a jump if $\left|\alpha_{+}-\alpha_{-}\right|\geq\delta_{0}$, indicating a (significant) discontinuity in function values, and as a kink if $\left|\alpha_{+}-\alpha_{-}\right|<\delta_{0}$ but $\left|\beta_{+}-\beta_{-}\right|\geq\delta_{0}$, indicating a (significant) discontinuity in slope without a (significant) jump in values, see Figure 1. The respective parameter spaces are denoted by $\Theta^{J}_{\delta_{0}}\coloneqq\left\{\theta\in\Theta_{\delta_{0}}:|\alpha_{+}-\alpha_{-}|\geq\delta_{0}\right\}$ for the jump case and $\Theta^{K}_{\delta_{0}}\coloneqq\Theta_{\delta_{0}}\setminus\Theta^{J}_{\delta_{0}}$ for the kink case.
@@
Formally, we define an online detector $\hat{\tau}_{n}$ as a Markov stopping time with respect to the filtration $\left\{\mathcal{F}_{t},1\leq t\leq n\right\}$, where $\mathcal{F}_{t}$ is the $\sigma$-algebra generated by observations $X_{1},\ldots,X_{t}$. Let $\mathcal{T}_{n}$ denote the collection of all such online detectors. Following Korostelev and Korosteleva, (2011, Chapter 6), we aim at achieving the convergence rate of the minimax quadratic risk of detection, as the sample size $n$ goes to infinity. Namely, we study the rate at which the quadratic risk
$R_{n}^{*}(\delta_{0})\;\coloneqq\;\inf_{\hat{\tau}_{n}\in\mathcal{T}_{n}}\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\right]$
converges to zero. This risk naturally balances type I (false detection) and type II (missed detection) errors, since it can be decomposed as
$\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\right]=\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\mathbb{I}\left(\hat{\tau}_{n}<\tau\right)\right]+\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\mathbb{I}\left(\hat{\tau}_{n}\geq\tau\right)\right].$ (2)
This quadratic risk framework, rooted in in-fill asymptotics, differs slightly from conventional frameworks in sequential change point detection (see Section 1.2 below), aligning more closely with offline change point analysis (see e.g. Frick et al., 2014a, ; Niu et al.,, 2016; Truong et al.,, 2020; Cho and Kirch,, 2024), where all observations are available upfront. By framing the online problem in a manner analogous to offline analysis, our approach facilitates a rigorous comparison between online and offline setups. Furthermore, the second term in (2), together with the false alarm probability, implies an upper bound on the expected detection delay as follows
$\mathbb{E}_{\tau}\left[\left(\hat{\tau}_{n}-\tau\right)^{2}\mathbb{I}\left(\hat{\tau}_{n}\geq\tau\right)\right]\geq\bigl{(}\mathbb{E}_{\tau}\left[\hat{\tau}_{n}-\tau\mid\hat{\tau}_{n}\geq\tau\right]\bigr{)}^{2}\bigl{(}1-\mathbb{P}_{\tau}(\hat{\tau}_{n}<\tau)\bigr{)},$ (3)
thereby connecting the quadratic risk to standard performance metrics in sequential change point detection (see Corollary 3). In addition, we put emphasis simultaneously on the statistical optimality and the computational and memory efficiency. Our contributions are threefold:
@@
    We introduce the Fast Limited-memory Optimal Change (FLOC) detector, which operates in constant time per observation and requires only constant memory, particularly independent of the sample size (see Proposition 4). The implementation of FLOC is provided in an R package available on GitHub under https://github.com/AnnikaHueselitz/FLOC.
@@
    We establish that the proposed FLOC detector achieves minimax optimal rates in quadratic risk for estimating the change point in segmented linear regression (see Theorems 1 and 5). Moreover, it reliably distinguishes between jump and kink types of structural changes with high probability (see Proposition 2).
  3. iii.
    Our theoretical analysis reveals a fundamental phase transition between jump and kink scenarios (see Table 1). Specifically, the minimax rate scales as $\log(n)/n$ for a jump and as $(\log(n)/n)^{1/3}$ for a kink, underscoring the different levels of detectability between abrupt and more gradual structural changes in the regression function.
@@
This phase transition between a jump and a kink has been similarly observed in offline change point estimation in segmented linear models (Chen,, 2021), see also Goldenshluger et al., (2006) and Shen et al., (2022) for general regression models. The literature on offline change point estimation in segmented linear models is notably extensive (see e.g. Bai and Perron,, 1998, Muggeo,, 2003, Frick et al., 2014b, , and more recently, Lee et al.,, 2016, Cho et al.,, 2025 and Cho and Li,, 2025 for high-dimensional extensions). In Table 1, we summarize the phase transition scenarios in segmented linear models, comparing online and offline setups in terms of optimal rates, computational efficiency and memory requirements. It is worth emphasizing that, unlike offline setups, where model validity is typically required globally over the whole domain of the function, our online detector FLOC is effective even with only local model validity around the change point. This flexibility makes our approach particularly well-suited for applications where model structures are reliable only within specific localized regions, as demonstrated with the real-world data set discussed in Section 4.4.
@@ 1.2 Related work
$\sup_{1\leq\kappa<\infty}\mathrm{D}_{\kappa}(\tilde{\tau})\qquad\text{subject to }\;\mathbb{E}\left[\tilde{\tau}\mid\text{no change}\right]\geq\gamma$
in the asymptotic regime of $\gamma\to\infty$. Here, $D_{\kappa}(\cdot)$ measures the detection delay when the change occurs at $\kappa$, and specific examples include
[…]
Within this framework, Yao, (1993) investigated the optimal detection of a jump, with respect to (4a), in linear regression models, and Yakir et al., (1999) studied the optimal detection of a kink with respect to (4b). One can relate such results to the model (1a)–(1c) in this paper by the correspondence of $\hat{\tau}=\tilde{\tau}/n$ and $n\approx\gamma$, see Corollary 3 for further details.
@@
As an addition to the existing literature, the proposed online detector FLOC achieves constant computational and memory complexity per observation, while also attaining minimax optimal rates in terms of quadratic risk for segmented linear regression models simultaneously for a jump and a kink (see again Table 1).
@@ 1.3 Organization and notation
The remainder of the paper is organized as follows. In Section 2, we introduce formaly the FLOC detector and establish its statistical guarantees in estimating the change point. The minimax lower bounds on convergence rates in quadratic risk are given in Section 3. Section 4 examines the empirical performance of the proposed FLOC detector on both simulated and real-world data sets and in comparison to state-of-the-art methods. All the proofs are deferred to Section 5, and Section 6 concludes the paper with a discussion.
Throughout, we denote by $\mathbb{P}_{\tau}\left(\cdot\right)$ the probability measure and by $\mathbb{E}_{\tau}\left[\cdot\right]$ the expectation under the model (1a)–(1c) with the true change point $\tau$. In particular, $\mathbb{P}_{0}\left(\cdot\right)$ and $\mathbb{E}_{0}\left[\cdot\right]$ refer to the case with no change (i.e. $\tau=0$). For sequences $\{a_{n}\}$ and $\{b_{n}\}$ of positive numbers, we use $a_{n}=\mathcal{O}(b_{n})$ to indicate that $a_{n}\leq Cb_{n}$ for some finite constant $C>0$, and write $a_{n}\asymp b_{n}$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$.
@@ 2 FLOC: Statistical theory and computational complexity
In this section, we introduce the Fast Limited-memory Optimal Change (FLOC) detector, based on two online detectors, one of which is used to detect a jump and the other to detect a kink. We then derive upper bounds on the quadratic risk of FLOC as well as its computational cost.
@@ 2.1 Risk bounds
Figure 2: Simulated scenarios illustrating a jump (upper panel) and a kink (lower panel). The black lines represent the underlying signal $f_{\theta}$, which exhibits a change at $\tau=516/n$, and the blue dots denote observations sampled from a Gaussian distribution with mean $f_{\theta}$, as specified in (1a). The red step functions show the values of the test statistic computed by FLOC. The vertical dotted lines indicate observation points where FLOC discards a bin; the bin sizes are $N_{J}=N_{K}=2$. The pre-change signal is estimated using the first $k=500$ observations. A change is flagged when the test statistic exceeds the threshold, with $\rho_{J}=1.5$ and $\rho_{K}=0.3$, shown as horizontal dashed lines. For both the jump and the kink scenarios, this occurs at observation $521$. For visual clarity in the lower panel (kink case), the threshold and test statistic values are scaled by a factor of 5.
@@
The core idea of the online detection algorithm is to estimate the pre-change signal $f_{-}$ at the beginning of the data sequence and then continuously assess whether the data within the most recent time window remains consistent with this underlying function. The online detector is built upon a certain test statistic, which reports a change if the test statistic exceeds a pre-specified threshold. To detect a jump, the test statistic is the mean of the residuals between the observed data and the estimated pre-change signal over the latest interval. For a kink, a weighted mean of these residuals is used instead, assigning greater weights to more recent observations.
@@
Consider a data sequence $X_{1},X_{2},\ldots,X_{n}$, and let $k,N_{J},N_{K}\leq n$ be natural numbers that may depend on $n$. A minimax rate optimal choice of these parameters will be provided in Theorem 1, while practical selection strategies are discussed later in Section 4.1. The function $f_{-}$ is estimated using the first $k$ data points via least squares, as follows:
@@
$\hat{f}_{-}\bigl{(}\frac{i}{n}\bigr{)}\coloneqq\hat{\alpha}+\hat{\beta}\frac{i}{n},\qquad\text{ with }\quad\bigl{(}\hat{\alpha},\hat{\beta}\bigr{)}\coloneqq\underset{\left(\alpha,\beta\right)\in\mathbb{R}^{2}}{\arg\min}\sum_{i=1}^{k}\left(\alpha+\beta\frac{i}{n}-X_{i}\right)^{2}.$ (5a)
We define the CUSUM jump test statistic of bin size $N_{J}$ as
$J_{m}\coloneqq\frac{1}{M_{J}}\sum_{i=1}^{M_{J}}\left(X_{m-M_{J}+i}-\hat{f}_{-}\bigl{(}\frac{m-M_{J}+i}{n}\bigr{)}\right),\qquad k<m\leq n,$ (5b)
@@
with $M_{J}\coloneqq 2N_{J}+\left(m\mod N_{J}\right)$, and similarly the CUSUM kink test statistic of bin size $N_{K}$ as
$K_{m}\coloneqq\frac{6}{M_{K}\left(M_{K}+1\right)\left(2M_{K}+1\right)}\sum_{i=1}^{M_{K}}i\left(X_{m-M_{K}+i}-\hat{f}_{-}\bigl{(}\frac{m-M_{K}+i}{n}\bigr{)}\right),\qquad k<m\leq n,$ (5c)
[…]
$\hat{\tau}_{J,n}\coloneqq\begin{cases}1&\text{if }|J_{m}|<\rho_{J}\text{ for all }m\in\{k+1,\ldots,n\},\\ \min\bigl{\{}m:|J_{m}|\geq\rho_{J},\ k<m\leq n\bigr{\}}/n&\text{otherwise,}\end{cases}$ (5d)
@@
$\hat{\tau}_{K,n}\coloneqq\begin{cases}1&\text{if }|K_{m}|<\rho_{K}\text{ for all }m\in\{k+1,\ldots,n\},\\ \min\bigl{\{}m:|K_{m}|\geq\rho_{K},\ k<m\leq n\bigr{\}}/n&\text{otherwise,}\end{cases}$ (5e)
@@
$\hat{\tau}_{n}\coloneqq\min\left(\hat{\tau}_{J,n},\hat{\tau}_{K,n}\right).$ (5f)
It is straightforward to verify that the detectors $\hat{\tau}_{J,n}$ and $\hat{\tau}_{K,n}$, as defined in (5d) and (5e), are Markov stopping times. As a consequence, FLOC, being the minimum of two Markov stopping times, is also a Markov stopping time, that is, FLOC is indeed an online detector (see also the proof in Section 5). An illustration of FLOC is provided in Figure 2. While the bin sizes $N_{J}$ and $N_{K}$ are fixed, the window sizes used by both the CUSUM jump and kink test statistics in (5b) and (5c) vary over time. This design offers an advantage in terms of storage efficiency: instead of retaining all recent observations and removing them individually, it suffices to store only the weighted sums of observations and discard entire bins when starting a new one (see Section 2.2).
@@
    Let $\hat{\tau}_{J,n}$ be the CUSUM jump detector in (5b) and (5d) with bin size $N_{J}=10^{3}\log(n)/(2c^{2})$ and threshold $\rho_{J}=4c/5$. Then, it holds for sufficiently large $n$,
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\right]\;\leq\;r^{*}_{J}<\infty,$
@@
    Let $\hat{\tau}_{K,n}$ be the CUSUM kink detector in (5c) and (5e) with bin size $N_{K}=(300/c^{2})^{1/3}n^{2/3}\log^{1/3}(n)$ and threshold $\rho_{K}=4c/(5n)$. Then, it holds for sufficiently large $n$,
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\hat{\tau}_{K,n}-\tau\right)\right)^{2}\right]\leq r^{*}_{K}<\infty,$
@@
    For the FLOC detector $\hat{\tau}_{n}$ defined in (5f), both of the above rates hold for sufficiently large $n$.